write an equation of the line that has the given line and passes through x-intercept and slope. x-intercept= 2, m= -2/3
Answer by ilana(307)

Since we know the x-intercept and the slope, we can write this equation in point-slope form, using the x-intercept 2 as the point (2,0).
So, we use the form: y-y1=m(x-x1), with m=slope, (x1,y1)=our point.
So we have y-0=-2/3(x-2), or y=-2/3x+4/3.